Shiw that the equations xp-yp=0, z(xp+yq)=2xy are compatible and solve them

1
Expert's answer
2021-12-30T17:43:40-0500

Solution:

Here, we may take "f=x p-y q, g=z(x p+y q)-2 x y" so that

"\dfrac{\partial(f, g)}{\partial(x, p)}=2 x y, \quad \dfrac{\partial(f, g)}{\partial(z, p)}=-x^{2} p, \quad \dfrac{\partial(f, g)}{\partial(y, q)}=-2 x y, \quad \dfrac{\partial(f, g)}{\partial(z, q)}=x y p"

from which it follows that

"[f, g]=x p(y q-x p)=0" (here [f,g] is the J, i.e. jacobian)

since "x p=y p" .

The equations are therefore compatible.

It is readily shown that "p=y \/ z, q=x \/ z" , so that we have to solve

which has solution

"z d z=y d x+x d y"

where "c_{1}" is a constant.

"z^{2}=c_{1}+2 x y"